OpenStudy (anonymous):

e^(4x) + 5e^(2x) - 24 = 0. How do you solve this type of problem?

OpenStudy (anonymous):

e^2x=y so e^(4x) + 5e^(2x) - 24 = 0=y^2+5y-24=(y+8)(y-3) y=3 ln3=2x x=ln(3)/2=0.55